OpenStudy (anonymous):

Megan bought 7 charms for $31.50. Each charms costs the same amount of money. Write an inequality that can be use to find the maximum amount of charms (c) Megan can buy with $75.00.

OpenStudy (anonymous):

4.5x<=75.0

OpenStudy (anonymous):

What is the maximum amount of charms Megan can buy with $75.00

OpenStudy (anonymous):

75/4.5

OpenStudy (anonymous):

I got 16.6666666667

OpenStudy (anonymous):

you can't buy .666666667 of a bracelet. so megan can buy up to 16

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

16 charms

OpenStudy (anonymous):

here is another one solve -2x-1 3x/2 +5/2 solve for x this is a grided response

OpenStudy (anonymous):

i dont understand the middle term. is it -13x/2

OpenStudy (anonymous):

no it is spouse to have an =

OpenStudy (anonymous):

-2x-1= 3x/2+ 5/2 -1= 3x/2 +2x +5/2 -1-5/2= 3x/2+2x -2/2-5/2= 3x/2+4x/2 -7/2= 7x/2 2(-7/2)= 7x -7/7=x x=-1

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

same one x+4.25=3.5x-1.5x-0.75

OpenStudy (anonymous):

2.555

OpenStudy (anonymous):

31.50/7=4.5 4.5* X= 75.00 75.00/4.5=16.6 x=16.6

OpenStudy (anonymous):

3/4x+2<_3x-1